Two different gasohol mixtures are available. One contains 12% ethanol and the other 5% ethanol. How much of each should be mixed to obtain 14 gal of ethanol having 10% alcohol?

Answer by ikleyn(52778)

.
Two different gasohol mixtures are available. One contains 12% ethanol and the other 5% ethanol.
How much of each should be mixed to obtain 14 gal of ethanol having 10% alcohol?
~~~~~~~~~~~~~~~~~~~

Mix x gallons of the 12% ethanol and (14-x) gallons of the 5% ethanol.


An equation for the amount of the pure ethanol in ingredient and the mixture is

    0.12x + 0.05*(14-x) = 0.1*14  gallons of ethanol.


Simplify and find x

    0.12x + 0.7 - 0.05x = 1.4

    0.12x - 0.05x = 1.4 - 0.7

        0.07x     =    0.7

            x     =   0.7/0.07 = 10.


ANSWER.  10 gallons of the 12% ethanol and 14-10 = 4 gallons of the 5% ethanol.

Solved.

----------------------

It is a standard and typical mixture problem.

For introductory lessons covering various types of mixture word problems see
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Typical word problems on mixtures from the archive
in this site.

You will find there ALL TYPICAL mixture problems with different methods of solutions,
explained at different levels of detalization,  from very detailed to very short.

Read them and become an expert in solution mixture word problems.